OpenStudy (anonymous):

Solve for B. A = 3(B + C) B = ((A-C) / (3)) B = ((3A-C) / (3)) B = ((A-3C) / (3))

OpenStudy (anonymous):

@Nnesha How do you make things a fraction on this website?

Nnesha (nnesha):

yep

Nnesha (nnesha):

http://prntscr.com/7fbjfw \frac

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

A

OpenStudy (anonymous):

the first choice? ""((A-C) / (3))""

Nnesha (nnesha):

``` \(\large\rm\frac{2}{2}\)

Nnesha (nnesha):

FIRST STEp!!! distribute parentheses by outside term

OpenStudy (anonymous):

3B + 3C

Nnesha (nnesha):

yes right A = 3B +3C move 3C to the left side and then solve for b

OpenStudy (anonymous):

3C + A = 3b \[\frac{ 3C }{ 3 } + \frac{ A }{ 3 } = \frac{ 3B }{ 3 }\]

Nnesha (nnesha):

how would you cancel out 3C frm the right side ?

OpenStudy (anonymous):